2 and half litres of oil are poured into a container whose cross section is a square of side 12 and half cm . How deep is the container?
Answer by Alan3354(69443)

2 and half litres of oil are poured into a container whose cross section is a square of side 12 and half cm . How deep is the container?
-----------
Vol = cross-sectional area * depth.
---
CSA = 12.5^2 = 156.25 sq cm
----
2.5 liters = 2500 cc (cubic cms)
----
depth of the oil = 2500/156.25
= 16 cm
-------------
The depth of the container is >= 16 cm.
It might be 16 cm.
It might be 20 cm.
It might be 500 cm.
------------------------
another math problem written by people who don't know what they're talking about.